When 0.617 g of sodium metal is added to an excess of hydrochloric acid, 6410 J of heat are produced. What is the enthalpy of the reaction as written?


2Na(s)+2HCl(aq)⟶2NaCl(aq)+H2(g)


1
Expert's answer
2021-04-25T07:35:09-0400

Mole of sodium=mass÷molar mass

Mole= 0.617÷23=0.026

∆H=q÷n

∆H= 6410÷0.026=246538.46J